# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 7 Acknowledgements
JFD specially thanks Hanzhong Liu, Julien Mairal, and Bin Yu for very useful clarifications regarding Ref. [MY12] and Lasso in general. JFD also thanks Yihui Quek for pointing out Ref. [QCR20] and Rahul Jain, Josep Lumbreras, and Marco Tomamichel for interesting discussions. DL acknowledges funding from the Latvian Quantum Initiative under EU Recovery and Resilience Facility under project no. 2.3.1.1.i.0/1/22/I/CFLA/001 in the final part of the project. CSP gratefully acknowledges Ministry of Education (MOE), AcRF Tier 2 grant (Reference No: MOE-T2EP20220-0013) for the funding of this research. This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05.
@@ Appendix A Summary of symbols
$\hat{\bm{\beta}}$ Lasso optimal solution
$\widetilde{\bm{\beta}}$ Lasso approximate solution
@@
$\alpha$ Mutual incoherence
$\gamma$ Mutual overlap

### v2

@@ 7 Acknowledgements
JFD thanks Hanzhong Liu and Bin Yu for very useful clarifications regarding Ref. [78] and Lasso in general. JFD thanks András Gilyén for Ref. [4], Yihui Quek for Ref. [85], and Rahul Jain, Josep Lumbreras, and Marco Tomamichel for interesting discussions. JFD also thanks Iosif Pinelis for answering a question on MathOverflow [69]. JFD acknowledges funding from ERC grant No. 810115-DYNASNET. DL acknowledges funding from the Latvian Quantum Initiative under EU Recovery and Resilience Facility under project no. 2.3.1.1.i.0/1/22/I/CFLA/001 in the final part of the project. CSP gratefully acknowledges Ministry of Education (MOE), AcRF Tier 2 grant (Reference No: MOE-T2EP20220-0013) for the funding of this research. This research is supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant and its Quantum Engineering Programme under grant NRF2021-QEP2-02-P05.
@@ Appendix A Summary of symbols
$\hat{\bm{\beta}}$ Optimal Lasso solution
$\widetilde{\bm{\beta}}$ Approximate Lasso solution
@@
$\alpha_{\mathcal{A}}$ Mutual incoherence
$\gamma_{\mathcal{A}}$ Mutual overlap